OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (anonymous):

@YanaSidlinskiy can you help me

OpenStudy (anonymous):

@chineseboy15

OpenStudy (anonymous):

Is the line for the equation y = - 8 horizontal or vertical? What is the slope of this line? Select the best answer. The line is horizontal. The slope is undefined. The line is horizontal. The slope is 0. The line is vertical. The slope is undefined. The line is vertical. The slope is 0.

OpenStudy (anonymous):

Select the coordinates of two points that are on the line x = 5. (5, 0) and (0, 5) (5, 0) and (5, -2) (5, 0) and (0, 2) (0, 5) and (0, 2)

OpenStudy (anonymous):

i need help with only those 3

OpenStudy (yanasidlinskiy):

The line x = 5 tells you that x MUST BE 5. It doesn't say anything about y so you get to pick anything for y that you please. Just make sure x is 5 as in (5,0) (5,3) (5,-90) and so on.

OpenStudy (yanasidlinskiy):

The slope y = -8 is horizontal and its slope is 0

OpenStudy (yanasidlinskiy):

Any questions? Any doubts??

OpenStudy (anonymous):

i dont get the one about the line=5

OpenStudy (anonymous):

would it be (5, 0) and (5, -2)

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (yanasidlinskiy):

Yes!:)

OpenStudy (anonymous):

Sorry it took me so long, i was stuck.

OpenStudy (yanasidlinskiy):

It's ok:) This website needs a server upgrade.

OpenStudy (yanasidlinskiy):

Thankx:)

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x - 7 y = - one halfx - 2 y = 2x - 2 y = - one halfx - 7 is my answer correct? i got A

OpenStudy (yanasidlinskiy):

Yup!:)